OpenStudy (anonymous):

Alex and Sam both left towns X and Y which are 280km apart at noon. Alex drives from X to Y at 60 km/h, whilst Sam drives from X to Y at 80km/h. At what time will they meet and where?

OpenStudy (cwrw238):

i assume you mean Sam drive for Y to X at 90 km/hr. for Alex dist x = 60 *t and for sam distance 280 -x = 80 *t so x = 60t and 280-x = 80 t solve this sytem of equations to find distance and time where they meet